two numbers b and c are chosen at random with replacement from the numbers 1,2,3,4,5,6,7,8,9 .find the probability that x2+bx+c>0 for all x belongs to RĀ 

Grade:11

3 Answers

deepak
7 years ago
if x2+bx+c>0
then it implies that discriminant should be less than zero
i.e. b2-4c>0
hence b>2\sqrt{c}
Ā 
if c=1 then b can be 3,4,5,6,7,8,9Ā  ā€¦..7
if c=2 then b can be 3,4,5,6,7,8,9Ā  ā€¦....7
if c=3 then b can be 4,5,6,7,8,9 ........6
if c=4 then b can be 5,6,7,8,9 ā€¦........5
if c=5 then b can be 5,6,7,8,9 ā€¦........5
if c=6 then b can be 5,6,7,8,9 ā€¦.ā€¦...5
if c=7 then b can be 6,7,8,9 ā€¦.ā€¦.4
if c=8 then b can be 6,7,8,9..........4
if c=9 then b can be 7,8,9 ā€¦.3
Ā 
therefore total no of favourable pairs are 46
Ā 
then the probability of the given even happening is 46/81 = 0.567
Ā 
hope it is useful
Ā 
deepak
7 years ago
sorry that is the incorrect solution this is the correct one
if x2+bx+c>0
then it implies that discriminant should be less than zeroĀ Ā 
i.e. b2-4cā€™
hence b\sqrt{c}
Ā 
if c=1 then b can be 1
if c=2 then b can be 1,2
if c=3 then b can be 1,2,3
if c=4 then b can be 1,2,3
if c=5 then b can be 1,2,3,4
if c=6 then b can be 1,2,3,4
if c=7 then b can be 1,2,3,4,5
if c=8 then b can be 1,2,3,4,5
if c=9 then b can be 1,2,3,4,5
Ā 
therefore total no of favourable pairs are 32
Ā 
then the probability of the given even happening is 32/81 = 0.395
Ā 
hope it was useful
deepak
7 years ago
sorry that is the incorrect solution this is the correct one
if x2+bx+c>0
then it implies that discriminant should be less than zeroĀ Ā 
i.e. b2-4c instead of
hence b\sqrt{c}
Ā 
if c=1 then b can be 1
if c=2 then b can be 1,2
if c=3 then b can be 1,2,3
if c=4 then b can be 1,2,3
if c=5 then b can be 1,2,3,4
if c=6 then b can be 1,2,3,4
if c=7 then b can be 1,2,3,4,5
if c=8 then b can be 1,2,3,4,5
if c=9 then b can be 1,2,3,4,5
Ā 
therefore total no of favourable pairs are 32
Ā 
then the probability of the given even happening is 32/81 = 0.395
Ā 
hope it was useful